OpenStudy (blackstreet23):

Find the radius and the interval of convergence sigma (X^k)/(k+1)

OpenStudy (amistre64):

first take the ratio, and work the limit

OpenStudy (blackstreet23):

\[\sum_{k=0}^{\infty} \frac{ X^k }{ k+1 }\]

OpenStudy (blackstreet23):

I did that well. I am having troubles testing the limits

OpenStudy (amistre64):

hmm, show your work, lets see how well you got to that point then

OpenStudy (blackstreet23):

I dont understand how to test if it decreasing in the alternating series test

OpenStudy (amistre64):

so your trying to test for x=-1 ?

OpenStudy (blackstreet23):

give me a second

OpenStudy (amistre64):

\[\lim_{k\to\infty}\frac{x^{n+1}/(k+2)}{x^n/(k+1)}\] \[\lim_{k\to\infty}x\frac{k+1}{k+2}\] \[|x|\lim_{k\to\infty}\frac{k+1}{k+2}=|x|\]

OpenStudy (amistre64):

when x=1, this is a harmonic series ... when x=-1, its an alternating harmonic series ... do you recall the convergence of those?

OpenStudy (blackstreet23):

harmonic no

OpenStudy (blackstreet23):

how do you do it?

OpenStudy (amistre64):

the harmonic does not converge, the alternating does ...

OpenStudy (blackstreet23):

but why? Could you explain?

OpenStudy (blackstreet23):

what is a harmonic series? and how do i know if it converges or diverges?

OpenStudy (amistre64):

harmonic is\[\frac11+\frac12+\frac13+\frac14+\frac15+...\] the proofing is textbook, but it eludes me

OpenStudy (blackstreet23):

so is like \[\frac{ 1 }{ K! }\]

OpenStudy (amistre64):

http://web.williams.edu/Mathematics/lg5/harmonic.pdf

OpenStudy (blackstreet23):

ohh my bad \[\frac{ 1 }{ k }\]

OpenStudy (blackstreet23):

ohh ok i see if it has a K on the bottom

OpenStudy (beginnersmind):

@amistre "the proofing is textbook, but it eludes me" 1/3 + 1/4 > 1/2 1/5 + 1/6 + 1/7 + 1/8 > 1/2 1/9 + ... 1/16 > 1/2 etc.

OpenStudy (blackstreet23):

I still dont get how can someone recognize that a function is a harmonic series just by looking at it

OpenStudy (amistre64):

the alternating harmonic series converges :) that proof eludes me as well ... just recall it from the classes.

OpenStudy (blackstreet23):

is because of the K?

OpenStudy (amistre64):

1/n is the harmonic series by definition

OpenStudy (amistre64):

yes

OpenStudy (blackstreet23):

so K by itself would be a harmonic series?

OpenStudy (blackstreet23):

2K +1

OpenStudy (amistre64):

1/k^p converges for p > 1

OpenStudy (blackstreet23):

10 k -2

OpenStudy (blackstreet23):

isnt it 1/k^p a p-series?

OpenStudy (amistre64):

yes

OpenStudy (blackstreet23):

ok sooo if it has a K is a harmonic series and if it has an exponent is a p series?

OpenStudy (amistre64):

1/k = 1/k^1 by definition, 1/k is the harmonic series.

OpenStudy (amistre64):

that is why the p series test says that if p > 1 it converges, but if 0<p<= 1 it diverges

OpenStudy (blackstreet23):

so it diverges?

OpenStudy (amistre64):

why do i feel like this is just going in circles ....

OpenStudy (blackstreet23):

and please special help in the alternating series one !

OpenStudy (blackstreet23):

sorry i am just reinforcing the knowledge you are giving me by enphasing what you say lol

OpenStudy (amistre64):

the alternating harmonic series converges, by properties that i cant recall either. they are explained better by the texts

OpenStudy (amistre64):

http://www.math.com/tables/expansion/tests.htm

OpenStudy (amistre64):

radius = 1 interval is: [-1,1)

OpenStudy (blackstreet23):

Ohh ok. I am so sleepy right now haha. I will analyse it more tomorrow :). Thanks a lot ! anything I will send you more messages :D

OpenStudy (amistre64):

im getting tired as well, but there are plenty of smarter people then me about :)